# Revisiting Convergence of AdaGrad with Relaxed Assumptions

### v1

@@ Abstract
In this study, we revisit the convergence of AdaGrad with momentum (covering AdaGrad as a special case) on non-convex smooth optimization problems. We consider a general noise model where the noise magnitude is controlled by the function value gap together with the gradient magnitude. This model encompasses a broad range of noises including bounded noise, sub-Gaussian noise, affine variance noise and the expected smoothness, and it has been shown to be more realistic in many practical applications. Our analysis yields a probabilistic convergence rate which, under the general noise, could reach at $\tilde{\mathcal{O}}(1/\sqrt{T})$. This rate does not rely on prior knowledge of problem-parameters and could accelerate to $\tilde{\mathcal{O}}(1/T)$ where $T$ denotes the total number iterations, when the noise parameters related to the function value gap and noise level are sufficiently small. The convergence rate thus matches the lower rate for stochastic first-order methods over non-convex smooth landscape up to logarithm terms [Arjevani et al., 2023]. We further derive a convergence bound for AdaGrad with mometum, considering the generalized smoothness where the local smoothness is controlled by a first-order function of the gradient norm.
@@ 1 Introduction
$\begin{split}\mathbb{E}\|g(\bm{x})-\nabla f(\bm{x})\|^{2}\leq B\|\nabla f(\bm{x})\|^{2}+C\quad\text{or}\quad\|g(\bm{x})-\nabla f(\bm{x})\|^{2}\leq B\|\nabla f(\bm{x})\|^{2}+C.\end{split}$ (1)
This noise model, verified in machine learning applications with feature noise [Fuller, 2009] and [Khani and Liang, 2020], and in robust linear regression [Xu et al., 2008], offers a more realistic portrayal by allowing the noise norm to increase with the gradient norm, covering both bounded noise and sub-Gaussian noise. These studies not only provided a convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$, but also addressed challenges posed by the entanglement of adaptive step-sizes and stochastic gradients, and the additional variance in (1). However, to the best of our knowledge, none of existing works have proved the convergence of vanilla AdaGrad under (1) without assuming bounded gradients. Moreover, the distinct step-size for each coordinate in AdaGrad, as opposed to an unified step-size for all coordinates in AdaGrad-Norm, brings more challenges when considering (1).
@@
where $f(\bm{x})\geq f^{*},\forall\bm{x}\in\mathbb{R}^{d}$. It’s obvious to verify that (2) is strictly weaker than (1), and thus than the bounded noise or sub-Gaussian noise condition [^1] For conciseness, we mainly consider the almost-sure version of this general noise model. Extending our high probability analysis from an almost-sure version to a sub-Gaussian version is easy, which will be included in Appendix.. Indeed, (2) could be regarded as an extension of (1) and the following expected smoothness condition [Gower et al., 2019, Grimmer, 2019, Wang and Yuan, 2023],
$\begin{split}\mathbb{E}\|g(\bm{x})\|^{2}\leq A(f(\bm{x})-f^{*})+B\quad\|g(\bm{x})\|^{2}\leq A(f(\bm{x})-f^{*})+B.\end{split}$ (3)
Existing researches have studied SGD’s convergence under (2) and smoothness, both in asymptotic [Poljak and Tsypkin, 1973] and non-asymptotic view [Khaled and Richtárik, 2023]. More importantly, it has been shown that numerous of practical stochastic gradient settings satisfy (2) but out of the range of (1), including commonly used perturbation, sub-sampling and compression [Khaled and Richtárik, 2023]. However, the analysis for SGD could not be directly extended to AdaGrad due to the correlation of adaptive-sizes and stochastic gradients, and the coordinate-wise performance in AdaGrad.
@@
This assumption was proposed by [Zhang et al., 2020b] through empirical studies on language models and later verified in large language models, e.g., [Zhang et al., 2020a, Crawshaw et al., 2022]. (4) generalizes the standard global smoothness condition and allows unbounded smooth parameter, bringing more challenges for the convergence analysis of adaptive methods. Previous works [Faw et al., 2023, Wang et al., 2023] have derived convergence bounds for AdaGrad-Norm with (1) and (4). Also, prior knowledge of problem-parameters is necessary as pointed out by the counter examples in [Wang et al., 2023]. However, the analysis for coordinate-wise AdaGrad is non-trivial and requires more delicate constructions, particular when considering the weaker noise assumption in (2).
In the following, we will summarize our main contributions.
@@
  • As direct corollaries, we also derive similar probabilistic convergence results of AdaGrad on non-convex smooth optimization with affine variance noise. More importantly, the convergence rate is optimal and adaptive to the noise level $C$ in (1).
@@ 2 Related works
Numerous of works mainly studied the convergence behavior of AdaGrad-Norm over non-convex smooth landscape. Li and Orabona [2019] first proved a convergence for AdaGrad-Norm. However, they studied a variant with a delayed step-size that is independent from the current stochastic gradient and required knowledge of the smoothness parameter for step-sizes. Getting rid of prior-knowledge on problem-parameters, Ward et al. [2020] relied on a novel proxy step-size technique and showed the convergence with an uniform bound of stochastic gradients for vanilla AdaGrad-Norm. Kavis et al. [2022] and Liu et al. [2023b] proved probabilistic convergence under the sub-Gaussian noise and the latter further abandoned the bounded gradient assumption. In case of the affine variance noise, Faw et al. [2022] proved a convergence while not requiring uniformly bounded gradients. However, their rate is adaptive to the noise level only when $B\sim\mathcal{O}(1/T)$. Wang et al. [2023] relied on a distinct framework to improve the dependency of the convergence rate with respect to $T$ and achieved the adaptivity on noise level without any restriction over $B$. Concurrently, Attia and Koren [2023] deduced a probabilistic bound, using a novel induction argument to control the function value gap. Their result also adapted to the noise level without further requirement on $B$. Liu et al. [2023a] formulated a convergence bound for AdaGrad-Norm and its acceleration version in quasar-convex smooth setting.
The element-wise version of AdaGrad was first studied in [Duchi et al., 2011] on online convex optimization. In non-convex smooth case, a line of works investigated AdaGrad with bounded gradients. Zou et al. [2019] explored the convergence of AdaGrad with a heavy-ball or Nesterov style momentum. Zhou et al. [2020] also covered AdaGrad in their analysis, but they deduced a bound under a relatively strict condition over gradients and involving a $1/\epsilon$ term in the bound. Défossez et al. [2022] studied AdaGrad with Adam-type momentum and improved the dependency on the momentum parameter $\beta$ to $\mathcal{O}((1-\beta)^{-1})$. Shen et al. [2023] introduced a weighted AdaGrad with unified momentum covering both heavy-ball and Nesterov’s acceleration. Recently, Liu et al. [2023b], a work mentioned before, derived a convergence bound for AdaGrad under coordinate-wise sub-Gaussian noise, i.e., $g(\bm{x})_{i}-\nabla f(\bm{x})_{i}$ is sub-Gaussian for each $i\in[d]$, without requiring bounded gradients.
@@
We briefly summarize some works on the convergence of SGD or AdaGrad with (1) and its variants under the non-convex smooth landscape. Bertsekas and Tsitsiklis [2000] provided an almost-sure convergence bound for SGD. In non-asymptotic view, Bottou et al. [2018] derived a convergence bound for SGD of the form $\mathcal{O}(1/T+\sqrt{C/T})$ when step-sizes are well tuned by the smooth parameter, $B$ and $C$. They also pointed out that the extension is immediate from the bounded noise case [Ghadimi and Lan, 2013]. The convergence of AdaGrad-Norm under (1) has been well studied by [Faw et al., 2022, Wang et al., 2023, Attia and Koren, 2023] as mentioned before. Faw et al. [2023] further extended the analysis considering a generalized smooth condition. However, none of these existing works could prove the convergence of coordinate-wise version of AdaGrad under (1) or a weaker noise condition.
@@
Since $\bm{x}^{*}$ is ill-defined for non-convex optimization, Gower et al. [2019] directly set $\mathbb{E}[\|g(\bm{x}^{*})\|^{2}]=B$ and deduced the non-convex version of the expected smoothness in (3), which aligns with the weak growth condition [Vaswani et al., 2019] when $B=0$. Gower et al. [2019] relied on (3) to analyze SGD over quasi-strongly convex optimization. Independently, Grimmer [2019] relied on (3) and developed a general framework for SGD equipped with projection operators over convex non-smooth functions. Wang and Yuan [2023] also used (3) to derive a convergence bound for SGD using bandwidth-based step-sizes.
@@
In conclusion, it’s clear to see that (2) is weaker than the above conditions including (1) (when assuming the existence of $f^{*}$), (3) and (5) [^2] We make the comparison when assuming all conditions are in almost-sure form.. Our result then shows that AdaGrad is powerful enough to converge to a stationary point under this mild noise assumption without prior knowledge of problem-parameters.
@@
Since its introduction by [Zhang et al., 2020b], the generalized smooth condition has been well studied under different algorithms, e.g., [Qian et al., 2021, Zhao et al., 2021, Reisizadeh et al., 2023]. Considering AdaGrad and its variants, Faw et al. [2023] established a convergence bound for AdaGrad-Norm considering (1). However, their result required $B<1$. Wang et al. [2023] further tightened the dependency to the iteration number $T$ and got rid of restriction on $B$.
@@ 3 Problem setting and algorithm
Given $\bm{x}\in\mathbb{R}^{d}$, we assume a gradient oracle that returns a random vector $g(\bm{x},\bm{z})\in\mathbb{R}^{d}$, where $\bm{z}$ denotes a random sample. The deterministic gradient of $f$ at $\bm{x}$ is denoted by $\nabla f(\bm{x})\in\mathbb{R}^{d}$.
@@
  • (A3) Relaxed affine variance noise: The gradient oracle satisfies that for some constants $A,B,C>0$, $\|g(\bm{x},\bm{z})-\nabla f(\bm{x})\|^{2}\leq A(f(\bm{x})-f^{*})+B\|\nabla f(\bm{x})\|^{2}+C,\forall\bm{x}\in\mathbb{R}^{d}$.
@@
for another three positive constants $A^{\prime},B^{\prime},C^{\prime}$. Therefore, (A3) is a generalization of (1) and (3). For more detailed examples of stochastic gradient settings satisfying Assumption (A3), we refer interested readers to see [Khaled and Richtárik, 2023, Proposition 2,3].
@@ 4 Main Convergence Result
$\frac{1}{T}\sum_{s=1}^{T}\|\nabla f(\bm{x}_{s})\|^{2}\leq\mathcal{O}\left[{\Delta}_{1}\left(\frac{B_{1}{\Delta}_{1}+\sqrt{B_{1}L\Delta}+\epsilon}{T}+\sqrt{\frac{A\Delta+C}{T}}\right)\right],$
@@
${\Delta}\sim$ $\mathcal{O}\left[f(\bm{x}_{1})-f^{*}+\frac{\sqrt{C}\eta d}{1-\beta}\log\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)+\frac{(A+B_{1}L)\eta^{2}d^{2}}{(1-\beta)^{3}}\log^{2}\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right].$
@@
With a simple calculation, when $\eta=c_{1}(1-\beta)^{3/2}$ for some constant $c_{1}>0$, the above upper bound has a minimum order of $\mathcal{O}((1-\beta)^{-1})$ with respect to $(1-\beta)^{-1}$.
@@
$\frac{1}{T}\sum_{s=1}^{T}\|\nabla f(\bm{x}_{s})\|^{2}\leq\mathcal{O}\left[{\Delta}_{1}\left(\frac{B_{1}{\Delta}_{1}+\sqrt{B_{1}L\Delta}+\epsilon}{T}+\sqrt{\frac{C}{T}}\right)\right],$
@@
${\Delta}\sim$ $\mathcal{O}\left[f(\bm{x}_{1})-f^{*}+{\sqrt{C}\eta d}\log\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)+{B_{1}L\eta^{2}d^{2}}\log^{2}\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right].$
@@
4) Assumption (A3) can be replaced by its sub-Gaussian form where
$\mathbb{E}_{\bm{z}}\left[\exp\left({\|g(\bm{x},\bm{z})-\nabla f(\bm{x})\|^{2}\over A(f(\bm{x})-f^{*})+B\|\nabla f(\bm{x})\|^{2}+C}\right)\right]\leq\mathrm{e},$ and our results still hold true, as shown in Appendix.

### v2

@@ Abstract
In this study, we revisit the convergence of AdaGrad with momentum (covering AdaGrad as a special case) on non-convex smooth optimization problems. We consider a general noise model where the noise magnitude is controlled by the function value gap together with the gradient magnitude. This model encompasses a broad range of noises including bounded noise, sub-Gaussian noise, affine variance noise and the expected smoothness, and it has been shown to be more realistic in many practical applications. Our analysis yields a probabilistic convergence rate which, under the general noise, could reach at $\tilde{\mathcal{O}}(1/\sqrt{T})$. This rate does not rely on prior knowledge of problem-parameters and could accelerate to $\tilde{\mathcal{O}}(1/T)$ where $T$ denotes the total number iterations, when the noise parameters related to the function value gap and noise level are sufficiently small. The convergence rate thus matches the lower rate for stochastic first-order methods over non-convex smooth landscape up to logarithm terms [Arjevani et al., 2023]. We further derive a convergence bound for AdaGrad with momentum, considering the generalized smoothness where the local smoothness is controlled by a first-order function of the gradient norm.
@@ 1 Introduction
$\begin{split}\mathbb{E}&\|g(\bm{x})-\nabla f(\bm{x})\|^{2}\leq B\|\nabla f(\bm{x})\|^{2}+C\\ \text{or}\quad&\|g(\bm{x})-\nabla f(\bm{x})\|^{2}\leq B\|\nabla f(\bm{x})\|^{2}+C.\end{split}$ (1)
This noise model, verified in machine learning applications with feature noise [Fuller, 2009, Khani and Liang, 2020], and in robust linear regression [Xu et al., 2008], offers a more realistic portrayal by allowing the noise norm to increase with the gradient norm, covering both bounded noise and sub-Gaussian noise. These studies not only provided a convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$, but also addressed challenges posed by the entanglement of adaptive step-sizes and stochastic gradients, and the additional variance in (1). However, to the best of our knowledge, none of existing works have proved the convergence of vanilla AdaGrad under (1) without assuming bounded gradients. Moreover, the distinct step-size for each coordinate in AdaGrad, as opposed to an unified step-size for all coordinates in AdaGrad-Norm, brings more challenges when considering (1).
@@
where $f(\bm{x})\geq f^{*},\forall\bm{x}\in\mathbb{R}^{d}$. It’s obvious to verify that (2) is strictly weaker than the almost surely affine variance noise in (1), and thus than the bounded noise or sub-Gaussian noise condition [^1] For conciseness, we mainly consider the almost-sure version of this general noise model. Extending our high probability analysis from an almost-sure version to a sub-Gaussian version is easy, which will be included in Appendix.. Indeed, (2) could be regarded as an extension of (1) and the following expected smoothness condition [Gower et al., 2019, Grimmer, 2019, Wang and Yuan, 2023],
$\begin{split}\mathbb{E}&\|g(\bm{x})\|^{2}\leq A(f(\bm{x})-f^{*})+B\\ \text{or}\quad&\|g(\bm{x})\|^{2}\leq A(f(\bm{x})-f^{*})+B.\end{split}$ (3)
Existing researches have studied SGD’s convergence behavior under (2) with smooth objective functions, both in asymptotic [Poljak and Tsypkin, 1973] and non-asymptotic view [Khaled and Richtárik, 2023]. More importantly, it has been shown that numerous of practical stochastic gradient settings satisfy (2) but out of the range of (1), including commonly used perturbation, sub-sampling and compression [Khaled and Richtárik, 2023]. However, the analysis for SGD could not be directly extended to AdaGrad due to the correlation of adaptive-sizes and stochastic gradients, and the coordinate-wise performance in AdaGrad.
@@
This assumption was proposed by [Zhang et al., 2020b] through empirical studies on language models and later verified in large language models, e.g., [Zhang et al., 2020a, Crawshaw et al., 2022]. (4) generalizes the standard global smoothness condition and allows unbounded smooth parameter, bringing more challenges for the convergence analysis of adaptive methods. Previous works [Faw et al., 2023, Wang et al., 2023] have derived convergence bounds for AdaGrad-Norm with (1) and (4). Also, prior knowledge of problem-parameters is necessary as pointed out by the counter examples in [Wang et al., 2023]. However, the analysis for coordinate-wise AdaGrad is non-trivial and requires more delicate constructions, particularly when considering the weaker noise assumption in (2).
In the following, we will summarize our main contributions as follows. We also refer readers to see the comparison of our results with existing works in Table 1 from the appendix.
@@
  • As direct corollaries, we also derive similar probabilistic convergence results of AdaGrad on non-convex smooth optimization with almost surely affine variance noise. More importantly, the convergence rate is optimal and adaptive to the noise level $C$ in (1).
@@ 2 Related works
Numerous of works mainly studied the convergence of AdaGrad-Norm over non-convex smooth landscape. Li and Orabona [2019] first proved the convergence for AdaGrad-Norm. However, they studied a variant with a delayed step-size that is independent from the current stochastic gradient and required knowledge of the smoothness parameter for tuning step-sizes. Getting rid of prior-knowledge on problem-parameters, Ward et al. [2020] relied on a novel proxy step-size technique and showed the convergence with an uniform bound of stochastic gradients for vanilla AdaGrad-Norm. Kavis et al. [2022] and Liu et al. [2023b] proved probabilistic convergence under the sub-Gaussian noise without relying on the bounded gradient assumption. In case of the affine variance noise, Faw et al. [2022] provided the convergence bound. However, their rate is adaptive to the noise level only when $B\sim\mathcal{O}(1/T)$. Wang et al. [2023] relied on a distinct framework to improve the dependency on $T$ in the convergece rate [Faw et al., 2022] and achieved the adaptivity on noise level without any restriction over $B$. Concurrently, Attia and Koren [2023] deduced a probabilistic bound, using a novel induction argument to control the function value gap. Their result also adapted to the noise level without further requirement on $B$. Liu et al. [2023a] formulated a convergence bound for AdaGrad-Norm and its acceleration version in quasar-convex smooth setting.
The element-wise version of AdaGrad was first studied in [Duchi et al., 2011] on online convex optimization. In non-convex smooth case, a line of works investigated AdaGrad with bounded gradients. Zou et al. [2019] explored the convergence of AdaGrad with a heavy-ball or Nesterov style momentum. Zhou et al. [2020] also covered AdaGrad in their analysis, but they deduced a bound under requiring bounded gradients’ summation. Défossez et al. [2022] studied AdaGrad with Adam-type momentum and improved the dependency on the momentum parameter $\beta$ to $\mathcal{O}((1-\beta)^{-1})$. Shen et al. [2023] introduced a weighted AdaGrad with unified momentum covering both heavy-ball and Nesterov’s acceleration. Recently, Liu et al. [2023b], a work mentioned before, derived a convergence bound under coordinate-wise sub-Gaussian noise, i.e., $g(\bm{x})_{i}-\nabla f(\bm{x})_{i}$ is sub-Gaussian for each $i\in[d]$, without requiring bounded gradients.
@@
We briefly summarize some works on the convergence of SGD or AdaGrad with (1) and its variants under the non-convex smooth landscape. Bertsekas and Tsitsiklis [2000] provided an almost-surely convergence bound for SGD. In non-asymptotic view, Bottou et al. [2018] derived a convergence bound for SGD of the form $\mathcal{O}(1/T+\sqrt{C/T})$ when step-sizes are well tuned by the smooth parameter, $B$ and $C$. They also pointed out that the extension is immediate from the bounded noise case [Ghadimi and Lan, 2013]. The convergence of AdaGrad-Norm under (1) has been well studied by [Faw et al., 2022, Wang et al., 2023, Attia and Koren, 2023] as mentioned before. Faw et al. [2023] further extended the analysis considering a generalized smooth condition. Recently, several studies have also examined the convergence of other adaptive methods, such as Adam [Hong and Lin, 2024], under (1). However, none of these existing works could prove the convergence of coordinate-wise version of AdaGrad under (1) or a weaker noise condition.
@@
Since $\bm{x}^{*}$ is ill-defined for non-convex optimization, Gower et al. [2019] then directly set $\mathbb{E}[\|g(\bm{x}^{*})\|^{2}]=B$ and deduced the non-convex version of the expected smoothness in (3), which aligns with the weak growth condition [Vaswani et al., 2019] when $B=0$. Gower et al. [2019] relied on (3) to analyze SGD over quasi-strongly convex optimization. Independently, Grimmer [2019] relied on (3) and developed a general framework for SGD equipped with projection operators over convex non-smooth functions. Wang and Yuan [2023] also used (3) to derive a convergence bound for SGD using bandwidth-based step-sizes.
@@
In conclusion, it’s clear to see that (2) is weaker than the above conditions including (1) (when assuming the existence of $f^{*}$), (3) and (5) [^2] We make the comparison when assuming all conditions are in almost-surely form.. Our result then shows that AdaGrad could find a stationary point under this mild noise assumption without prior knowledge of problem-parameters.
@@
The generalized smooth condition [Zhang et al., 2020b] has been well studied under different algorithms, e.g., [Qian et al., 2021, Zhao et al., 2021, Reisizadeh et al., 2023, Zhang et al., 2020a, Crawshaw et al., 2022]. Considering AdaGrad and its variants, Faw et al. [2023] established a convergence bound for AdaGrad-Norm considering (1). However, their result required $B<1$. Wang et al. [2023] further tightened the dependency to the iteration number $T$ and got rid of restriction on $B$.
@@ 3 Problem setting and algorithm
Given $\bm{x}\in\mathbb{R}^{d}$, we assume a gradient oracle that returns a random vector $g(\bm{x},\bm{z})\in\mathbb{R}^{d}$, where $\bm{z}$ denotes a random sample. The gradient of $f$ at $\bm{x}$ is denoted by $\nabla f(\bm{x})\in\mathbb{R}^{d}$.
@@
  • (A3) Relaxed almost surely affine variance noise: The gradient oracle satisfies that for some constants $A,B,C>0$, $\|g(\bm{x},\bm{z})-\nabla f(\bm{x})\|^{2}\leq A(f(\bm{x})-f^{*})+B\|\nabla f(\bm{x})\|^{2}+C,a.s.,\forall\bm{x}\in\mathbb{R}^{d}$.
@@
for another three positive constants $A^{\prime},B^{\prime},C^{\prime}$. Therefore, (A3) is a generalization of the almost surely noise models in (1) and (3). For more detailed examples of stochastic gradient settings satisfying Assumption (A3), we refer interested readers to see [Khaled and Richtárik, 2023, Proposition 2,3].
@@ 4 Main Convergence Result
$\frac{1}{T}\sum_{s=1}^{T}\|\nabla f(\bm{x}_{s})\|^{2}$
$\leq$ $\mathcal{O}\left[{\Delta}_{1}\left(\frac{B_{1}{\Delta}_{1}+\sqrt{B_{1}L\Delta}+\epsilon}{T}+\sqrt{\frac{A\Delta+C}{T}}\right)\right],$
@@
${\Delta}\sim$ $\mathcal{O}\left[f(\bm{x}_{1})-f^{*}+\frac{\sqrt{C}\eta d}{1-\beta}\log\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right.$
$\left.+\frac{(A+B_{1}L)\eta^{2}d^{2}}{(1-\beta)^{3}}\log^{2}\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right].$
@@
With a simple calculation, when $\eta=c_{1}(1-\beta)^{3/2}$ for some constant $c_{1}>0$, the above upper bound has a minimum order of $\mathcal{O}((1-\beta)^{-1})$ with respect to $(1-\beta)^{-1}$. The comparison of existing results with our convergence bound could be found in Table 1 from the appendix.
@@
$\frac{1}{T}\sum_{s=1}^{T}\|\nabla f(\bm{x}_{s})\|^{2}$
$\leq$ $\mathcal{O}\left[{\Delta}_{1}\left(\frac{B_{1}{\Delta}_{1}+\sqrt{B_{1}L\Delta}+\epsilon}{T}+\sqrt{\frac{C}{T}}\right)\right],$
@@
${\Delta}\sim$ $\mathcal{O}\left[f(\bm{x}_{1})-f^{*}+{\sqrt{C}\eta d}\log\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right.$
$\left.+{B_{1}L\eta^{2}d^{2}}\log^{2}\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right].$
@@
4) Assumption (A3) can be replaced by its sub-Gaussian form where $\mathbb{E}_{\bm{z}}\left[\exp\left({\|g(\bm{x},\bm{z})-\nabla f(\bm{x})\|^{2}\over A(f(\bm{x})-f^{*})+B\|\nabla f(\bm{x})\|^{2}+C}\right)\right]\leq\mathrm{e},$ and our results still hold true, as shown in Appendix.